OpenStudy (anonymous):

I need to prove that log5(3) is not a rational number. I'm not quite sure of how to prove this...

OpenStudy (anonymous):

I think that for log5(3) to be rational, lg(3)/lg(5) must also be a rational number

OpenStudy (shubhamsrg):

well,let log5(3) = a then 5^a = 3 simplifies the thinking now hmm ?

OpenStudy (hba):

a=1

OpenStudy (anonymous):

wel, if that is proof enough then it was quite easy

OpenStudy (shubhamsrg):

i dont think still its been proven or it'll be worthwhile carrying forward this way.. lets hear others.. @hba what do you mean by a=1 ?

OpenStudy (hba):

I wanted to play a let game but it wont work here. But we can see, A rational number is a number which can be written as a fraction of two integers, where the denominator is nonzero. (Another way of saying the same thing: A rational number is a number that can be written in the form p/q, where p/q are integers and q ≠ 0.) So it would be a=log(3)+2i pi n/log(5) Where n is a set of integers.

OpenStudy (hba):

I do not know how can i prove it is not a rational number though :/

hartnn (hartnn):

continuing what @shubhamsrg have done well,let log5(3) = a ..... (where a is a rational number) then 5^a = 3 since, a is rational number, it can be expressed in the form a=b/c where, b and c are INTEGERS. so, we have. \(\huge \sqrt[c]{5^b}=3 \implies5^b=3^c\) now there are no pair of integers b,c which can satisfy this ! hence a cannot be rational number. this is proof by contradiction.